### v1

# Common Inpainted Objects In-N-Out of Context

Abstract
We present Common Inpainted Objects In-N-Out of Context (COinCO), a novel dataset addressing the scarcity of out-of-context examples in existing vision datasets. By systematically replacing objects in COCO images through diffusion-based inpainting, we create 97,722 unique images featuring both contextually coherent and inconsistent scenes, enabling effective context learning. Each inpainted object is meticulously verified and categorized as in- or out-of-context through a multimodal large language model assessment. Our analysis reveals significant patterns in semantic priors that influence inpainting success across object categories. We demonstrate three key tasks enabled by COinCO: (1) training context classifiers that effectively determine whether existing objects belong in their context; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique levels, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning. COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision and image forensics. Our code and data are at: https://github.com/YangTianze009/COinCO.
1 Introduction
Context is fundamental to visual understanding [1, 2, 3]. When humans view a scene, we instinctively assess the contextual coherence between objects and their environment. This context-based reasoning is essential for interpreting real-world scenes [4] and beyond. Take a look at Figure 2—can you spot the inpainted (i.e., fake) object in each image? Contextual understanding helps identify objects that violate scene expectations, even when pixel-level artifacts are imperceptible. By using context, humans naturally assess whether objects appear in plausible settings—a potted plant flying in the sky would immediately raise suspicion, while the same plant in a garden appears perfectly natural.
Learning context from data, however, is difficult. A significant challenge is that unusual scenes with out-of-context objects are rare in real life. Common computer vision datasets [5, 6, 7, 8] primarily contain objects in their natural settings, creating a scarcity of examples with contextual violations. This presents a fundamental obstacle: how can we train machine learning models that require large data to recognize contextual inconsistencies when such examples are inherently uncommon?
To address this challenge, we propose a dataset with both in-context and out-of-context objects through controlled image manipulation. By systematically replacing objects in existing real scenes, we can generate the necessary examples of contextual violations while preserving the overall structure of the scene. We specifically chose the Common Objects in Context (COCO) dataset [5] as our foundation because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
Figure 1: Which object is fake? Only one object per image is inpainted. Out-of-context inpainted objects are easier to identify. Answers are revealed at the bottom of this page [^2] Figure 2 answers. 1st row: bird, laptop; 2nd row: hotdog, potted plant, potted plant; 3rd row: horse, horse..
Using Stable Diffusion’s inpainting model [9], we replace exactly one object per COCO image. This selective approach allows us to maintain the broader scene context while introducing precise, controlled variations in object-scene relationships. Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Multimodal Large Language Model (MLLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
Figure 2 shows the pipeline of our data creation and the downstream tasks. We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), reflecting our emphasis on context for inpainted objects. Our dataset contains 97,722 unique inpainted images, each accompanied by rich annotations.
COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model). COinCO serves as a challenging benchmark, encouraging new research for the computer vision community.
The main contributions of our work are:
1. We introduce a novel, large-scale dataset of partially manipulated images that enhances COCO, featuring strategically inpainted objects that are either contextually coherent or inconsistent with their scenes.
2. We employ COinCO’s context labels to train a context classifier. This task utilizes the visual features of an image, alongside the semantic features of detected objects to predict an object to be in- or out-of-context.
3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level categories for possible objects that fit the given context.
4. We demonstrate that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art methods without fine-tuning.
2 Related Work
COCO and its extensions. The Common Objects in Context (COCO) dataset [5] is a widely used benchmark for various computer vision tasks, including object detection, instance segmentation, and captioning. Over the years, several extensions have been proposed. Some datasets primarily focus on expanding annotations without modifying image content. COCO-WholeBody refines human keypoint detection by adding facial, hand, and foot keypoints [10], while LVIS introduces a long-tailed distribution with over 1,000 categories for instance segmentation [7]. COCO-Stuff incorporates background (stuff) annotations for panoptic segmentation [6], and RefCOCO enables referring expression-based object localization [8]. Other works improve vision-language alignment, such as COCO-Caption for image captioning [11] and COCO-Text for scene text detection [12]. Additionally, CD-COCO, which applies image distortions to test robustness [13], and COCONut, which unifies multiple segmentation tasks [14], explore scene complexity. However, these datasets do not manipulate contextual relationships. Our dataset uniquely reconstructs scene context by replacing original objects with out-of-context alternatives via inpainting, creating a testbed for context modeling and fake localization that enables the learning of complex scene semantics.
Context reasoning. Context is essential for understanding object relationships and finding anomalies in complex scenes. Biederman et al. [4] identified five relational principles—support, interposition, probability, position, and size—that help identify contextual inconsistencies. Prior works apply these principles using object co-occurrence and support relationships for out-of-context detection [2]. Others have demonstrated that slight context changes in “object transplanting” experiments can cause significant errors in object detection. Acharya et al. [15] introduced a Graph Contextual Reasoning Network (GCRN) to model co-occurrence and relative position to detect out-of-context objects. A comprehensive review of context in vision is in [16]. Ours is the first work to feature a mixture of in- and out-of-context objects in large-scale data. Moreover, context is under-explored in fake detection. Multimodal Large Language Models (MLLMs)  [17, 18, 19, 20, 21], offer a human-aligned approach to nuanced context reasoning. We are the first to use context in fake detection with MLLMs.
Figure 2: Our COinCO pipeline. (a) For a given COCO image, an object is randomly replaced by with Stable Diffusion inpainting. (b) Inpainting success is verified using object detection and an MLLM. Successes are added to the dataset, while fail cases are regenerated and retested. (c) Inpainted images are classified as in-context or out-of-context using the MLLM. (d) Instance-level (object category) and clique-level (supercategory) information is predicted from the inpainted image using an MLP and recovered with Stable Diffusion. (e) Fake pixels are detected with pretrained fake detectors and enhanced within the bounding box region of a detected out-of-context object. (f) Semantic and visual embeddings are derived from the object list and inpainted image, respectively. These embeddings are used to train an MLP for classifying instances as in- or out-of-context.
Fake image generation and detection. Image manipulation has a long history from traditional object insertion techniques [22, 23] to GAN-based synthesis [24, 25, 26, 27, 28]. Recently, diffusion models set new standards [29, 30], with text-to-image models like Stable Diffusion [9] and DALL-E [31] ControlNet enhances diffusion models with controls [32]. Emerging techniques tailor diffusion methods for object manipulation, inpainting, and harmonization [33, 34], 3D [35], and relighting [36, 37].
For detecting manipulated images, earlier works classify at image-level [38, 39]. Advanced methods detect various manipulations with details: PSCC-Net [40] employs spatio-channel correlation across scales, CAT-Net [41] focuses on JPEG compression artifacts, and ManTra-Net [42] captures diverse manipulation traces. TruFor combines RGB and noise-sensitive fingerprints to produce anomaly and confidence maps [43]. For fake image datasets, e.g., GenImage [44], CIFAKE [45], DE-FAKE [30], they focus on fully-synthesized images. Diffusion-based inpainting datasets like COCOGlide [43] and TGIF [46] utilize COCO images but they have limited image quantity, object replacements constrained to the same category (thus lacking out-of-context objects), and no context labels. Our dataset contains significant out-of-context fake objects in real scenes with rich characteristics.
3 Building COinCO
We propose COinCO, a novel context-oriented inpainted objects dataset derived from the widely-used Common Objects in Context (COCO) Dataset [5]. Leveraging COCO’s image- and object-level annotations, COinCO enhances COCO by systematically replacing objects with diffusion-based inpainting, providing images with enriched contextual diversity. This rendition has 95,320 unique training images and 2,402 testing images from the COCO2017 training and validation sets, respectively. Each inpainted object is categorized as in- or out-of-context and is accompanied by information about the original object and its replacement.
For each COCO image, we (1) randomly pick an object from the image and perform inpainting using Stable Diffusion, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with an MLLM to classify the new object as in- or out-of-context.
3.1 Inpainting
(a) Inpainting success rate.
(b) Context reasoning prompt.
Figure 3: (a) Inpainting success rate for original-inpainted object pairs. Rows are classes of original objects and columns are inpainted objects. A darker color indicates higher success rate. (b) Context reasoning prompt for Molmo [17].
We use Stable Diffusion 2 [47] for inpainting. Inpainting a new object requires an original image, an inpainting mask, and a prompt. For each COCO image, we randomly select an object. The object’s mask is slightly dilated and enclosed in a bounding box, which is used as the inpainting mask. In the preliminary study, we observed that the use of dilation and bounding boxes reduced residual artifacts from the original objects. For the replacement object, we randomly select one of COCO’s 80 object categories and use it as the text prompt.
The vanilla Stable Diffusion inpainting pipeline often struggles to inpaint small objects. To overcome this, we crop the original image around an enlarged inpainting area, inpaint, and then scale the results back to the original size. We use alpha blending to seamlessly merge the inpainted region with the original image. For more details on inpainting, see the supplementary materials.
3.2 Automated inpainting verification
Diffusion-based inpainting is imperfect: there is no guarantee the target object will be successfully inpainted. To confirm the success of inpainting, we use a COCO-trained YOLOv8x [48] object detector. If the inpainted object is detected within the inpainting mask, the process is deemed successful. For failed cases, we complete two additional rounds of inpainting and verification, pairing YOLOv8x with a state-of-the-art MLLM (Molmo [17]) to capture YOLO’s false negatives. Images that fail inpainting for all three rounds are discarded, resulting in 97,722 successful images.
Interestingly, we found that inpainting success rate is highly correlated with semantic priors. Figure 3(a) shows the inpainting success proportions for all original-inpainted object pairs. We note dense neighborhoods of inpainted object classes with increased probability of inpainting success. Using average linkage hierarchical clustering, we grouped replacement objects based on inpainting success proportions. With GloVe embeddings, we calculated cosine similarity between items within and across clusters. Mann-Whitney U tests revealed a significantly higher sample mean similarity within clusters (0.169) vs. across clusters (0.119) (bootstrapped p =.$=.002$). The cluster with the highest inpainting success included nearly all animals and food indicating a bias for these classes, while lowest performing cluster included furniture and household objects, which are highly context-dependent. A Kruskal-Wallis test confirmed significant differences across COCO supercategories (p $<0.001$), with food and animals outperforming electronics, appliances, and furniture. These findings validate the semantic coherence of the clusters formed and the vertical block pattern noted in Figure 3(a). Additional statistics are in the supplementary material. To our knowledge, this is the first thorough analysis of semantic priors in diffusion-based inpainting tasks.
3.3 Context reasoning
We label context for all inpainted objects. An object is considered in-context if it adheres to the contextual coherence of the scene with existing objects, while violations of the coherence result in an out-of-context classification. Our context reasoning is based on three fundamental principles regarding context  [2]: location, size, and co-occurrence. Location examines the spatial positioning of objects within a scene. Size evaluates whether the object’s dimensions are proportionate to the scene geometry. Co-occurrence measures whether objects’ simultaneous presence is common. To label context accordingly, we leverage the power of the Molmo [17]. We design a deliberate prompt to guide Molmo in performing language and visual reasoning for the context classification of our inpainted images (see Figure 3(b)). Provided with an inpainted image and context criteria, Molmo can conduct holistic visual-semantic reasoning that a pure language-based model cannot.
3.4 Comparison with prior datasets
Table 1: COinCO significantly enriches COCO with inpainting and context labels.
COCO (2017) COCO-Stuff COCOGlide COinCO
Dataset size 123K 164K 512 97,722
Unique source images 123K 164K 512 97,722
Replacement classes - - same as original 80
Out-of-context images - - -
Context reasoning - -
Inpainting - -
Labels objects objects, scenes (fake) objects (fake) objects, context
COinCO extends COCO by introducing contextual diversity with inpainted objects (Table 1). Unlike COCO and COCO-Stuff, which primarily focus on object and scene (stuff) annotations, COinCO provides intentional out-of-context scenarios, enabling research in contextual reasoning and fake localization. While COCOGlide also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
4 Manual Verification
To validate the reliability of our automated models for object detection and context classification, we conducted a comprehensive manual verification study. We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context, based on the aforementioned principles.
4.1 Verification on inpainted object detection
We first evaluated whether target objects can be reliably detected in the inpainted region. We compared COCO-trained detectors, including YOLO [48], MMDetection [49], and GroundingDINO [50] on 1,000 human-labeled images in Table 2. AR: Alignment Rate, PR: Precision, RE: Recall, FPR: False Positive Rate. AR is the primary metric reflecting the agreement between a model and a human.
Method AR (%) PR RE FPR
YOLOv8x 76.4 93.78 67.86 8.15
YOLOv11x 66.6 94.80 50.93 5.06
MMDet 67.6 86.87 58.54 16.01
GroundDINO 59.5 96.14 38.66 2.81
Table 2: Object detection performance.
Model Acc (%) Prec (%) Rec (%) F1 (%)
VisualNet 52.98 51.82 84.62 64.28
SemanticNet 76.33 78.19 73.02 75.52
VisSemanticNet 76.90 77.66 75.51 76.57
Table 3: Context classification performance.
For ground truth, annotators recorded whether the inpainted object was successfully integrated within the mask. We tuned object detector parameters, such as confidence, IoU threshold, and augmentation. YOLOv8x achieved the highest alignment rate (AR) in relation to the manual labels (76.4%)–AR is the rate at which the model agrees with human annotators. While GroundingDINO displayed marginally higher precision (96.14%) and a lower false positive rate (2.81%), the model’s recall was much lower than YOLOv8x (38.66% versus 67.86%), reflecting its failure to identify positives.
In addition, we tested Molmo [17] using cropped inpainting mask regions as input, prompting it to confirm whether the target object was present. Molmo achieved an AR of $82.6\%$ with human annotations, surpassing the performance of conventional object detectors. Thus, YOLOv8x and Molmo work together in our cascaded verification process for efficiency and accuracy.
4.2 Verification on context classification
To verify context classification, we established strict criteria. First, an agreement was required between Molmo and annotators regarding the presence of the target object (as in Section 4.1). Second, all human annotators were to agree on the context classification (in- or out-of-context). Of our 1000 manually annotated images, a total of 477 instances met these criteria. Among these instances, Molmo’s context classification aligned with human annotations in 370 cases ($77.57\%$). These results confirm Molmo’s reliability in context understanding and, together with YOLOv8x, validate our cascaded detection and classification verification process.
5 Analysis on COinCO
With COinCO as our testbed, we explore the role of context in several tasks. We demonstrate how our data can be used to learn a general model to classify objects as in- or out-of-context. We then propose a novel Objects-from-Context task for predicting which objects naturally belong in a given scene. Finally, we evaluate state-of-the-art fake detection models and show how contextual information can enhance fake localization without any fine-tuning. These applications demonstrate the versatility of COinCO and underscore the fundamental role of context in visual understanding. Examples of our pipeline are illustrated in Figure 4.
Original Mask Inpainted Fake Det. Context Reasoning Objects-from-Context
Refer to caption Refer to caption Refer to caption Refer to caption “A kite can reasonably be flying in the air above a railway track… The kite appears to be appropriately sized for flying, neither too large nor too small for the environment.
Final decision: In-context.
Refer to caption Refer to caption Refer to caption Refer to caption A cow should not be in the sky. It is impossible for a cow to appear together with the moon.
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A horse in a bathroom, especially next to urinals, is highly unusual and unreasonable… Horses and urinals are completely unrelated objects that should not appear together…
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption ”Cups and water are common together, but a cup floating on its own in a natural body of water is not reasonable.”
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A soccer ball is reasonable in a bedroom setting…Soccer balls can appear together with beds, as they are both common items in a bedroom.
Final decision: In-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A refrigerator is highly unusual and out of place in a zoo enclosure, especially one with giraffes and other animals…
Final decision: Out-of-context.”
Figure 4: Inpainting, fake detection, and objects-from-context results. Context reasoning responses are color-coded by location, size, and co-occurrence. Original objects are in red. Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator, orange.
5.1 In- and out-of-context classification
COinCO’s context labels enables the training of a generalized context model to classify objects as in- or out-of-context within an image, addressing the question: “Does this object belong in this context?" We train a binary classification model that outputs a single score through a sigmoid activation function, where $1$: “out-of-context” and 00: “in-context.” For training and evaluation, we create a balanced dataset by supplementing our generated samples with original COCO images. As objects in COCO are naturally occurring, we assume they are “in-context”, ensuring a 50:50 ratio. We evaluate three models with different input modalities: (1) VisualNet: Uses only visual features from the image and mask via the VAE encoder in Stable Diffusion. (2) SemanticNet: Uses only semantic embeddings from BERT [51], including the average embedding of existing objects and the embedding of the query object. (3) VisSemanticNet: Uses both visual and semantic embeddings.
According to Table 3, the significant performance gap between VisualNet and SemanticNet shows that contextual understanding is primarily a reasoning task. VisualNet struggles to capture the complex relationships and real-world knowledge required for context reasoning and tends to predict objects as out-of-context, resulting in high recall but low accuracy and precision. VisSemanticNet achieves the best overall performance. Despite their simplicity, our context classification models show promising results and we hope they will inspire further research for more advanced context-aware models.
5.2 Objects-from-Context prediction
While our context classification model determines if an existing object belongs in a scene, our Objects-from-Context prediction task tackles a different question: “What object(s) fit to this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [52]. We formulate this task at two levels: 1) Instance-level prediction. Predicting the exact object class among COCO’s 80 classes. 2) Clique-level prediction. Predicting the clique an object belongs to, using COCO’s super-categories: accessory, animal, appliance, electronic, food, furniture, indoor, kitchen, outdoor, person, sports, and vehicle.
We design a model that takes two inputs: the inpainted image and a binary mask indicating the target region. During training, we use the dilated bounding box of the original object as the mask and its class as the label. We use the VAE encoder from Stable Diffusion to extract latents from both inputs, which are processed by an MLP that predicts across COCO’s 80 classes. For clique-level evaluation, we map predicted object classes to their corresponding super-categories and check if they match the original objects. Importantly, our model can handle arbitrary mask sizes and locations during inference, making it adaptable for broader context reasoning applications. Detailed model architectures and training information are in the supplementary material. For comparison, we implement a baseline that ranks candidate objects based on the co-occurrence frequency of other objects in a scene [53]. If no other objects are present, this baseline defaults to random selection.
Table 4: Instance-level and clique-level object-from-context prediction accuracy.
Method Instance-level (%) Clique-level (%)
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
Random 1.25 3.75 6.25 8.33 25.00 41.67
Co-occurrence [53] 1.54 4.70 7.29 9.37 30.72 52.91
Ours 16.32 31.89 42.80 35.10 61.41 78.31
Table 4 shows our model’s superior grasp of contextual features, significantly outperforming random and co-occurrence baselines. At the instance level, the substantial improvement in top-3 and top-5 accuracy (31.89% and 42.80%) indicates that even when the model’s top prediction is incorrect, the true object is often ranked among the most probable candidates. Our model demonstrates stronger performance at the clique level, correctly predicting semantically similar objects. This result demonstrates that our approach effectively learns the relationship between scene contexts and the objects that belong within them. Our model offers three key applications as shown in the three rows of Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [54]; (2) suggesting contextually appropriate objects based on spatial location within a scene, enabling intelligent image editing [33]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation. This versatility advances context-driven visual understanding across multiple practical applications.
Refer to caption Refer to caption Refer to caption
Instance P(%) skis: 31.6 snowboard: 24.3 kite: 14.8 surfboard: 7.0 backpack: 5.9 Clique P(%) sports: 79.2 accessory: 6.6 person: 4.9 vehicle: 4.2 animal: 3.0 Instance P(%) kite: 22.9 skis: 22.1 snowboard: 19.6 surfboard: 8.6 person: 5.5 Clique P(%) sports: 75.1 accessory: 6.1 vehicle: 5.7 person: 5.5 animal: 4.2 Instance P(%) horse: 5.0 zebra: 5.0 giraffe: 4.9 person: 4.6 elephant: 4.4 Clique P(%) animal: 37.2 vehicle: 14.9 sports: 12.1 outdoor: 10.7 accessory: 6.8
Figure 5: Object-from-Context prediction. A red box is a query. The top row shows three examples (two inpainted, one original COCO), and the bottom row lists instance-level and clique-level predictions with their probabilities (P). Objects in red are the top predictions.
Method Acc F1 AUC AP
ManTraNet [42] 85.4 30.7 84.9 50.7
Trufor [43] 89.7 55.9 93.4 73.6
PSCC-Net [40] 89.5 46.8 95.4 79.5
CAT-Net [41] 92.7 76.5 97.4 90.3
Table 5: SOTA fake detection performance.
Figure 6: Context enhancement results.
5.3 Context-empowered fake localization
Fake localization aims to identify the specific synthetic regions in an image, producing a pixel-level map that is fine-grained compared to binary image-level classification. In COinCO, the ground truth mask for fake localization is the original object’s entire bounding box used in inpainting. We evaluate localization performance using four common pixel-level metrics [41, 40]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
State-of-the-art performance. We benchmark several SOTA fake detection models on COinCO (Table 5). The high accuracy and AUC scores ($>84\%$ for all models) can be attributed to the small average size of inpainted regions in our data. As most images are majority authentic, these metrics are biased towards high values. F1 and AP, which better reflect the precision of fake localization, reveal a clear performance gap among models. CAT-Net achieves the best performance across all metrics with $92.65\%$ accuracy, $76.47\%$ F1 score, $97.35\%$ AUC, and $90.33\%$ AP. Trufor obtains the second-best F1 score of $55.92\%$, while PSCC-Net shows strong performance in AUC ($95.38\%$) and AP ($79.52\%$). ManTraNet demonstrates relatively lower performance, particularly in F1 score ($30.68\%$) and AP ($50.72\%$). These findings highlight significant differences and room for improvement among SOTA fake localizers. Examples of these SOTA models’ predictions are shown in Figure 7.
Incorporating context in fake localization. To leverage contextual information for improved fake localization, we propose a simple yet effective way of enhancing the prediction scores of fake localization models in regions corresponding to out-of-context objects. Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; otherwise $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
Image GT [41] [42] [43] [40]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Fake localization performance.
We evaluate this context-enhancement approach under two settings. In the oracle setting, we use ground truth annotations to enhance predictions within the fake object’s mask region when it is out-of-context. This serves as an upper bound for performance gains. In the practical setting, the fake objects are unknown. We propose the use of Molmo to identify suspicious objects based on size and location, deliberately excluding co-occurrence to prevent false positives. For instance, if an image contains only two objects, e.g., an apple and an inpainted traffic light, either might be seen as out-of-context. By focusing on physical violations, our method remains robust while effectively detecting fake objects. As Molmo may identify multiple out-of-context objects in an image, we define $M_{\text{OC}}$ as the union of these objects’ masks. In both settings, we set the enhancement factor $\gamma$ to $5$ to enhance prediction scores in out-of-context regions. A detailed analysis of $\gamma$ is in supplementary materials.
Figure 6 shows how our context-enhancement improves fake localization for all SOTA detectors without any finetuning. The effectiveness of our method stems from the complementary nature of context detection and fake localization. This combination ensures that even if Molmo misidentifies authentic objects as out-of-context, enhancement only strengthens the base model’s predictions, preserving robustness. These results highlight context as a valuable signal for fake detection, opening new possibilities for context-aware image forensics.
6 Conclusion and Limitations
We present COinCO, a novel dataset designed to feature in- and out-of-context objects in real scenes. By strategically replacing objects in COCO with diffusion-based inpainting, we systematically diversify the contextual status of objects in complex scenes. COinCO’s context labels allow us to train a general context classifier. We also introduce a novel Object-from-Context prediction task. Finally, our work advances current fake detection approaches by leveraging contextual information during fake pixel localization. COinCO and our findings highlight the importance of context in visual understanding and provide a foundation and testbed for future research.
While COinCO advances context-aware visual understanding, context classification remains inherently subjective despite our structured criteria, as human annotators and MLLMs occasionally disagree on contextual coherence. The Objects-from-Context prediction task is currently limited to COCO’s predefined categories. Future work could explore extending these tasks to open-vocabulary settings for more flexible context reasoning.

### v2

@@
 By systematically replacing objects in COCO images through diffusion-based inpainting, we create 97,722 unique images featuring both contextually coherent and inconsistent scenes, enabling effective context learning.
-Each inpainted object is meticulously verified and categorized as in- or out-of-context through a multimodal large language model assessment.
-Our analysis reveals significant patterns in semantic priors that influence inpainting success across object categories.
-We demonstrate three key tasks enabled by COinCO: (1) training context classifiers that effectively determine whether existing objects belong in their context; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique levels, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning.
-COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision and image forensics.
-Our code and data are at: https://github.com/YangTianze009/COinCO.
+Each inpainted object is meticulously verified and categorized as in- or out-of-context through Large Vision Language Model assessments.
+We demonstrate three key tasks enabled by COinCO: (1) a fine-grained context reasoning approach that classifies objects as in- or out-of-context based on three criteria; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique level semantics, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning.
+COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision, including image forensics.
+Code and dataset are available at https://co-in-co.github.io/.
 1 Introduction
-Context is fundamental to visual understanding [1, 2, 3].
-When humans view a scene, we instinctively assess the contextual coherence between objects and their environment.
-This context-based reasoning is essential for interpreting real-world scenes [4] and beyond.
-Take a look at Figure 2—can you spot the inpainted (i.e., fake) object in each image?
-Contextual understanding helps identify objects that violate scene expectations, even when pixel-level artifacts are imperceptible.
-By using context, humans naturally assess whether objects appear in plausible settings—a potted plant flying in the sky would immediately raise suspicion, while the same plant in a garden appears perfectly natural.
+Context is fundamental to visual understanding [3, 43, 45, 66, 13, 61].
+When humans view a scene, we instinctively assess the coherence between objects and their environment.
+This context-based reasoning is essential for judging whether objects naturally belong in a scene [5].
+Take a look at Figure 1—do the red-boxed objects seem reasonable in their contexts?
+Context reasoning can reveal out-of-place objects, providing a complementary cue for detecting manipulated content.
+By using context, humans assess whether objects appear in plausible settings—a horse in a grassy field with trees in the background appears perfectly natural, while a tiny zebra on a beach next to a person immediately raises suspicion due to its unusual size and location.
 Learning context from data, however, is difficult.
 A significant challenge is that unusual scenes with out-of-context objects are rare in real life.
-Common computer vision datasets [5, 6, 7, 8] primarily contain objects in their natural settings, creating a scarcity of examples with contextual violations.
+Common computer vision datasets [34, 10, 21, 31] primarily contain objects in their natural settings, resulting in a scarcity of examples with contextual violations.
 This presents a fundamental obstacle: how can we train machine learning models that require large data to recognize contextual inconsistencies when such examples are inherently uncommon?
-To address this challenge, we propose a dataset with both in-context and out-of-context objects through controlled image manipulation.
-By systematically replacing objects in existing real scenes, we can generate the necessary examples of contextual violations while preserving the overall structure of the scene.
-We specifically chose the Common Objects in Context (COCO) dataset [5] as our foundation because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
-Figure 1: Which object is fake?
-Only one object per image is inpainted.
-Out-of-context inpainted objects are easier to identify.
-Answers are revealed at the bottom of this page [^2] Figure 2 answers.
-1st row: bird, laptop; 2nd row: hotdog, potted plant, potted plant; 3rd row: horse, horse..
-Using Stable Diffusion’s inpainting model [9], we replace exactly one object per COCO image.
+Figure 1: COinCO contains a rich set of inpainted out-of-context objects (top two rows) and in-context objects (bottom row).
+To address this challenge, we propose a new dataset with both in-context and out-of-context objects through controlled image manipulation.
+By systematically replacing objects in existing real scenes, we generate the necessary examples of contextual violations while preserving the overall structure of the scene.
+We specifically build on top of the Common Objects in Context (COCO) dataset [34] because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
+Using a diffusion-based inpainting model [48], we replace exactly one object per COCO image.
 This selective approach allows us to maintain the broader scene context while introducing precise, controlled variations in object-scene relationships.
-Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Multimodal Large Language Model (MLLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
+Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Large Vision Language Model (LVLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
+We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), emphasis on context for inpainted objects.
+It contains 97,722 inpainted images, each accompanied by rich annotations.
 Figure 2 shows the pipeline of our data creation and the downstream tasks.
-We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), reflecting our emphasis on context for inpainted objects.
-Our dataset contains 97,722 unique inpainted images, each accompanied by rich annotations.
-COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model).
-COinCO serves as a challenging benchmark, encouraging new research for the computer vision community.
-The main contributions of our work are:
+COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including fine-grained context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model).
+Our main contributions are:
 1.
@@
 2.
-We employ COinCO’s context labels to train a context classifier.
-This task utilizes the visual features of an image, alongside the semantic features of detected objects to predict an object to be in- or out-of-context.
-3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level categories for possible objects that fit the given context.
-4. We demonstrate that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art methods without fine-tuning.
+We introduce a fine-grained context reasoning task in COinCO based on three criterion-specific dimensions and train efficient student models through distillation from a 72B LVLM teacher.
+These models provide modular, interpretable, and high-accuracy context reasoning for inpainted and real images.
+3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level semantic categories for possible objects that fit the given context.
+4. We show that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art image forensics methods without fine-tuning.
+Figure 2: COinCO pipeline.
+Left: Building COinCO.
+We start with COCO images and annotations.
+For each image, we randomly select an object and replace it using Stable Diffusion inpainting.
+We verify the inpainted object using object detection and a vision-language model.
+Successfully inpainted images are added to the dataset, while failed cases are regenerated and retested.
+Finally, each inpainted object is labeled as in-context or out-of-context based on three criteria: location, size, and co-occurrence.
+Right: COinCO Tasks.
+We demonstrate three downstream applications enabled by COinCO.
+Sec 5.1: Fine-grained context classification evaluates objects on each criterion using vision-language models.
+Sec 5.2: Objects-from-context prediction identifies which object categories fit a given context at both instance and clique level semantics, and can recover the original object with an image generation model.
+Sec 5.3: Context-enhanced fake detection improves pretrained fake detectors by enhancing predictions in out-of-context regions.
 2 Related Work
@@
 Context reasoning.
-Context is essential for understanding object relationships and finding anomalies in complex scenes.
+Context is essential for understanding object relationships and finding anomalies in complex scenes [43].
 Biederman et al.
-[4] identified five relational principles—support, interposition, probability, position, and size—that help identify contextual inconsistencies.
-Prior works apply these principles using object co-occurrence and support relationships for out-of-context detection [2].
-Others have demonstrated that slight context changes in “object transplanting” experiments can cause significant errors in object detection.
-Acharya et al.
-[15] introduced a Graph Contextual Reasoning Network (GCRN) to model co-occurrence and relative position to detect out-of-context objects.
-A comprehensive review of context in vision is in [16].
+[5] identified relational principles such as support, probability, and size that help identify contextual inconsistencies, which prior works apply using co-occurrence and support relationships for out-of-context detection [13].
+Others have shown that slight context changes can cause significant errors in object detection, and Acharya et al.
+[1] introduced a Graph Contextual Reasoning Network modeling co-occurrence and relative position to detect out-of-context objects.
+Context is particularly important for object detection [17, 7]; a comprehensive review is in [58].
 Ours is the first work to feature a mixture of in- and out-of-context objects in large-scale data.
-Moreover, context is under-explored in fake detection.
-Multimodal Large Language Models (MLLMs)  [17, 18, 19, 20, 21], offer a human-aligned approach to nuanced context reasoning.
-We are the first to use context in fake detection with MLLMs.
-Figure 2: Our COinCO pipeline.
-(a) For a given COCO image, an object is randomly replaced by with Stable Diffusion inpainting.
-(b) Inpainting success is verified using object detection and an MLLM.
-Successes are added to the dataset, while fail cases are regenerated and retested.
-(c) Inpainted images are classified as in-context or out-of-context using the MLLM.
-(d) Instance-level (object category) and clique-level (supercategory) information is predicted from the inpainted image using an MLP and recovered with Stable Diffusion.
-(e) Fake pixels are detected with pretrained fake detectors and enhanced within the bounding box region of a detected out-of-context object.
-(f) Semantic and visual embeddings are derived from the object list and inpainted image, respectively.
-These embeddings are used to train an MLP for classifying instances as in- or out-of-context.
+While context remains under-explored in fake detection, Large Vision Language Models (LVLMs) [15, 36, 35, 54, 28] offer a human-aligned approach to context reasoning; we are the first to leverage them for context-based fake detection.
 Fake image generation and detection.
-Image manipulation has a long history from traditional object insertion techniques [22, 23] to GAN-based synthesis [24, 25, 26, 27, 28].
-Recently, diffusion models set new standards [29, 30], with text-to-image models like Stable Diffusion [9] and DALL-E [31] ControlNet enhances diffusion models with controls [32].
-Emerging techniques tailor diffusion methods for object manipulation, inpainting, and harmonization [33, 34], 3D [35], and relighting [36, 37].
-For detecting manipulated images, earlier works classify at image-level [38, 39].
-Advanced methods detect various manipulations with details: PSCC-Net [40] employs spatio-channel correlation across scales, CAT-Net [41] focuses on JPEG compression artifacts, and ManTra-Net [42] captures diverse manipulation traces.
-TruFor combines RGB and noise-sensitive fingerprints to produce anomaly and confidence maps [43].
+Image manipulation has evolved from traditional object insertion [18, 9] to GAN-based synthesis [19, 25, 8, 29, 30] and diffusion models [23, 49], including Stable Diffusion [48], DALL-E [46], and ControlNet [65].
+Emerging techniques further tailor diffusion methods for object manipulation, inpainting, and harmonization [11, 64, 52], 3D [37], and relighting [22, 26].
+For detection, earlier works classify at image-level [14, 42], while advanced methods localize manipulations: PSCC-Net [39] uses spatio-channel correlation, CAT-Net [32] exploits JPEG artifacts, ManTra-Net [60] captures manipulation traces, and TruFor [20] combines RGB and noise-sensitive fingerprints.
 For fake image datasets, e.g., GenImage [67], CIFAKE [6], DE-FAKE [49], they focus on fully-synthesized images.
@@
 We propose COinCO, a novel context-oriented inpainted objects dataset derived from the widely-used Common Objects in Context (COCO) Dataset [34].
+We specifically choose COCO as our foundation for several reasons.
+First, we focus on common, everyday objects where contextual reasoning is most meaningful and widely applicable.
+COCO’s 80 classes span diverse supercategories covering frequent real-world categories such as animals, food, vehicles, and household items, making it the standard benchmark for context-rich scenes.
+Second, COCO provides reliable instance masks and annotations essential for high-quality inpainting.
+Expanding to rare or unusual objects would dilute the contextual focus and introduce annotation noise, as contextual expectations become less defined for uncommon categories.
 Leveraging COCO’s image- and object-level annotations, COinCO enhances COCO by systematically replacing objects with diffusion-based inpainting, providing images with enriched contextual diversity.
-This rendition has 95,320 unique training images and 2,402 testing images from the COCO2017 training and validation sets, respectively.
-Each inpainted object is categorized as in- or out-of-context and is accompanied by information about the original object and its replacement.
-For each COCO image, we (1) randomly pick an object from the image and perform inpainting using Stable Diffusion, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with an MLLM to classify the new object as in- or out-of-context.
+Our dataset comprises 97,722 successfully inpainted images derived from COCO2017, with each inpainted object categorized as in- or out-of-context with reasoning and accompanied by information about the original object and its replacement.
+Task-specific evaluation protocols are detailed in their respective sections.
+For each COCO image, we (1) randomly pick an object from the image and inpaint a new object, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with LVLMs to classify the new object as in- or out-of-context.
 3.1 Inpainting
-(a) Inpainting success rate.
-(b) Context reasoning prompt.
-Figure 3: (a) Inpainting success rate for original-inpainted object pairs.
-Rows are classes of original objects and columns are inpainted objects.
-A darker color indicates higher success rate.
-(b) Context reasoning prompt for Molmo [17].
-We use Stable Diffusion 2 [47] for inpainting.
+We use Stable Diffusion [48] for inpainting.
 Inpainting a new object requires an original image, an inpainting mask, and a prompt.
 For each COCO image, we randomly select an object.
-The object’s mask is slightly dilated and enclosed in a bounding box, which is used as the inpainting mask.
-In the preliminary study, we observed that the use of dilation and bounding boxes reduced residual artifacts from the original objects.
-For the replacement object, we randomly select one of COCO’s 80 object categories and use it as the text prompt.
-The vanilla Stable Diffusion inpainting pipeline often struggles to inpaint small objects.
+The object’s mask is slightly dilated and enclosed in a bounding box as the inpainting mask.
+Dilation and bounding boxes reduce residual artifacts from the original objects.
+For replacement, we randomly select one of COCO’s 80 categories as the text prompt.
+The vanilla inpainting pipeline often struggles to inpaint small objects.
 To overcome this, we crop the original image around an enlarged inpainting area, inpaint, and then scale the results back to the original size.
 We use alpha blending to seamlessly merge the inpainted region with the original image.
-For more details on inpainting, see the supplementary materials.
-3.2 Automated inpainting verification
-Diffusion-based inpainting is imperfect: there is no guarantee the target object will be successfully inpainted.
-To confirm the success of inpainting, we use a COCO-trained YOLOv8x [48] object detector.
-If the inpainted object is detected within the inpainting mask, the process is deemed successful.
-For failed cases, we complete two additional rounds of inpainting and verification, pairing YOLOv8x with a state-of-the-art MLLM (Molmo [17]) to capture YOLO’s false negatives.
-Images that fail inpainting for all three rounds are discarded, resulting in 97,722 successful images.
-Interestingly, we found that inpainting success rate is highly correlated with semantic priors.
-Figure 3(a) shows the inpainting success proportions for all original-inpainted object pairs.
-We note dense neighborhoods of inpainted object classes with increased probability of inpainting success.
-Using average linkage hierarchical clustering, we grouped replacement objects based on inpainting success proportions.
-With GloVe embeddings, we calculated cosine similarity between items within and across clusters.
-Mann-Whitney U tests revealed a significantly higher sample mean similarity within clusters (0.169) vs.
-across clusters (0.119) (bootstrapped p =.$=.002$).
-The cluster with the highest inpainting success included nearly all animals and food indicating a bias for these classes, while lowest performing cluster included furniture and household objects, which are highly context-dependent.
-A Kruskal-Wallis test confirmed significant differences across COCO supercategories (p $<0.001$), with food and animals outperforming electronics, appliances, and furniture.
-These findings validate the semantic coherence of the clusters formed and the vertical block pattern noted in Figure 3(a).
-Additional statistics are in the supplementary material.
-To our knowledge, this is the first thorough analysis of semantic priors in diffusion-based inpainting tasks.
+We found that for stronger inpainting models, such as Flux [33] and Qwen Image [59], they focus more on visual fidelity on in-context objects and tend to sacrifice the ability to generate out-of-context objects and provide less precise spatial control over the inpainting region.
+More details on inpainting and the analysis of different inpainting models are in the supplementary materials.
+Figure 3: Context reasoning prompt for LVLMs.
 3.2 Context reasoning
-We label context for all inpainted objects.
-An object is considered in-context if it adheres to the contextual coherence of the scene with existing objects, while violations of the coherence result in an out-of-context classification.
-Our context reasoning is based on three fundamental principles regarding context  [2]: location, size, and co-occurrence.
-Location examines the spatial positioning of objects within a scene.
-Size evaluates whether the object’s dimensions are proportionate to the scene geometry.
-Co-occurrence measures whether objects’ simultaneous presence is common.
-To label context accordingly, we leverage the power of the Molmo [17].
-We design a deliberate prompt to guide Molmo in performing language and visual reasoning for the context classification of our inpainted images (see Figure 3(b)).
-Provided with an inpainted image and context criteria, Molmo can conduct holistic visual-semantic reasoning that a pure language-based model cannot.
+We annotate context labels to all inpainted objects through a multi-model consensus process.
+Rather than relying on one single large model, three state-of-the-art LVLMs (Molmo-72B [15], Qwen2.5-VL-72B [2], and InternVL3.5-38B [57]) independently evaluate each inpainted image according to three context criteria: location, size, and co-occurrence.
+We design a structured prompt (Figure 3) to guide the LVLMs in performing this contextual reasoning.
+For every image, each model first provides criterion-level reasoning by examining (1) whether the spatial placement of the object is appropriate for the scene layout, (2) whether its scale aligns with the scene geometry, and (3) whether its simultaneous presence with existing objects is contextually plausible.
+An object is labeled as out-of-context if it violates any of these criteria; otherwise, it is deemed in-context.
+After generating criterion-level analyses, each LVLM produces a final binary judgment (in-context vs.
+out-of-context).
+To ensure the reliability of our context annotations, we retain labels only for images where a unanimous agreement is reached.
+This consensus requirement yields 73,929 images for which we release both the final in-/out-of-context labels and the complete criterion-level reasoning from all LVLMs.
+The remaining images are still part of COinCO, but we do not provide context annotations for them due to labeling ambiguity.
 3.3 Comparison with prior datasets
-Table 1: COinCO significantly enriches COCO with inpainting and context labels.
-COCO (2017) COCO-Stuff COCOGlide COinCO
+Table 1: COinCO significantly enhances COCO in new ways.
+COCO (2017) COCO-Stuff COCOGlide Ours
 Dataset size 123K 164K 512 97,722
@@
 Inpainting - -
-Labels objects objects, scenes (fake) objects (fake) objects, context
+Labels objects objects, scenes (fake) objects (fake) objects, context reasoning
 COinCO extends COCO by introducing contextual diversity with inpainted objects (Table 1).
 Unlike COCO and COCO-Stuff, which primarily focus on object and scene (stuff) annotations, COinCO provides intentional out-of-context scenarios, enabling research in contextual reasoning and fake localization.
-While COCOGlide also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
+While COCOGlide [20] also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
 4 Manual Verification
 To validate the reliability of our automated models for object detection and context classification, we conducted a comprehensive manual verification study.
-We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context, based on the aforementioned principles.
+We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context from human judgment.
 4.1 Verification on inpainted object detection
-We first evaluated whether target objects can be reliably detected in the inpainted region.
-We compared COCO-trained detectors, including YOLO [48], MMDetection [49], and GroundingDINO [50] on 1,000 human-labeled images in Table 2.
-AR: Alignment Rate, PR: Precision, RE: Recall, FPR: False Positive Rate.
-AR is the primary metric reflecting the agreement between a model and a human.
-Method AR (%) PR RE FPR
-YOLOv8x 76.4 93.78 67.86 8.15
-YOLOv11x 66.6 94.80 50.93 5.06
-MMDet 67.6 86.87 58.54 16.01
-GroundDINO 59.5 96.14 38.66 2.81
-Table 2: Object detection performance.
-Model Acc (%) Prec (%) Rec (%) F1 (%)
-VisualNet 52.98 51.82 84.62 64.28
-SemanticNet 76.33 78.19 73.02 75.52
-VisSemanticNet 76.90 77.66 75.51 76.57
-Table 3: Context classification performance.
-For ground truth, annotators recorded whether the inpainted object was successfully integrated within the mask.
-We tuned object detector parameters, such as confidence, IoU threshold, and augmentation.
-YOLOv8x achieved the highest alignment rate (AR) in relation to the manual labels (76.4%)–AR is the rate at which the model agrees with human annotators.
-While GroundingDINO displayed marginally higher precision (96.14%) and a lower false positive rate (2.81%), the model’s recall was much lower than YOLOv8x (38.66% versus 67.86%), reflecting its failure to identify positives.
-In addition, we tested Molmo [17] using cropped inpainting mask regions as input, prompting it to confirm whether the target object was present.
-Molmo achieved an AR of $82.6\%$ with human annotations, surpassing the performance of conventional object detectors.
-Thus, YOLOv8x and Molmo work together in our cascaded verification process for efficiency and accuracy.
+To select an appropriate detector for our automated inpainting verification pipeline, we evaluated several COCO-trained models on 1,000 human-labeled images where annotators recorded whether the target object was successfully inpainted within the mask region.
+As shown in Table 2, while GroundingDINO [38] achieved the highest precision (96.14%), its low recall (38.66%) would discard too many successful inpainting results.
+YOLOv8x [47] achieved the best balance with an F1 score of 78.75%, combining high precision (93.78%) with reasonable recall (67.86%).
+We therefore selected YOLOv8x as our detector and use it to verify inpainted objects: if YOLOv8x detects the target object within the inpainting mask, the process is deemed successful; for failed cases, we complete up to two additional rounds of inpainting and verification; images that fail all three rounds are discarded, resulting in 97,722 successful images.
+Notably, false positives often correspond to lower-quality inpaintings where the target object is present but poorly rendered, rather than complete inpainting failures.
+Table 2: Object detection performance on inpainted objects.
+Metric YOLOv8x YOLOv11x MMDet GroundDINO
+Precision 93.78 94.80 86.87 96.14
+Recall 67.86 50.93 58.54 38.66
+F1 78.75 66.26 69.95 55.14
 4.2 Verification on context classification
-To verify context classification, we established strict criteria.
-First, an agreement was required between Molmo and annotators regarding the presence of the target object (as in Section 4.1).
-Second, all human annotators were to agree on the context classification (in- or out-of-context).
-Of our 1000 manually annotated images, a total of 477 instances met these criteria.
-Among these instances, Molmo’s context classification aligned with human annotations in 370 cases ($77.57\%$).
-These results confirm Molmo’s reliability in context understanding and, together with YOLOv8x, validate our cascaded detection and classification verification process.
-5 Analysis on COinCO
+Following the procedure in Section 3.2, we retained 73,929 images where all three LVLMs reached unanimous agreement on the final decision, with 64,879 labeled as out-of-context and 9,050 as in-context.
+To validate the quality of these context annotations, we randomly sampled 1,000 images (300 in-context, 700 out-of-context) for human verification.
+Human annotators classified each image using the same context criteria.
+The LVLMs consensus achieved 95.10% agreement with human annotations, confirming the high reliability of our labels.
+Among the 1,000 annotated images, we identified 164 cases where the three LVLMs agreed on the final out-of-context decision but disagreed on at least one specific criterion.
+Human evaluation of these disagreements revealed that Qwen2.5VL-72B achieved the highest accuracy at 87.80% (144/164).
+For the remaining 836 images where all LVLMs agreed on both the final decision and all criteria, human annotators assessed reasoning quality.
+Qwen2.5VL-72B received 94.70% preference for its detailed and interpretable explanations.
+Based on these validation results, we adopt Qwen2.5VL-72B’s reasoning as the primary annotations for all 73,929 images in our dataset, with the consensus from MolMo and InternVL serving to ensure annotation reliability through multi-model verification.
+5 COinCO-enabled Context Tasks
 With COinCO as our testbed, we explore the role of context in several tasks.
-We demonstrate how our data can be used to learn a general model to classify objects as in- or out-of-context.
+We demonstrate how our data can be used to train efficient context classifiers through knowledge distillation, fine-tuning smaller models to evaluate objects based on our three context criteria.
 We then propose a novel Objects-from-Context task for predicting which objects naturally belong in a given scene.
-Finally, we evaluate state-of-the-art fake detection models and show how contextual information can enhance fake localization without any fine-tuning.
+Finally, we evaluate state-of-the-art fake detection models and show that context can enhance fake localization without any fine-tuning.
 These applications demonstrate the versatility of COinCO and underscore the fundamental role of context in visual understanding.
-Examples of our pipeline are illustrated in Figure 4.
+Examples of those tasks are in Figure 4.
 Original Mask Inpainted Fake Det. Context Reasoning Objects-from-Context
@@
 Original objects are in red.
-Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator, orange.
-5.1 In- and out-of-context classification
-COinCO’s context labels enables the training of a generalized context model to classify objects as in- or out-of-context within an image, addressing the question: “Does this object belong in this context?" We train a binary classification model that outputs a single score through a sigmoid activation function, where $1$: “out-of-context” and 00: “in-context.” For training and evaluation, we create a balanced dataset by supplementing our generated samples with original COCO images.
-As objects in COCO are naturally occurring, we assume they are “in-context”, ensuring a 50:50 ratio.
-We evaluate three models with different input modalities: (1) VisualNet: Uses only visual features from the image and mask via the VAE encoder in Stable Diffusion.
-(2) SemanticNet: Uses only semantic embeddings from BERT [51], including the average embedding of existing objects and the embedding of the query object.
-(3) VisSemanticNet: Uses both visual and semantic embeddings.
-According to Table 3, the significant performance gap between VisualNet and SemanticNet shows that contextual understanding is primarily a reasoning task.
-VisualNet struggles to capture the complex relationships and real-world knowledge required for context reasoning and tends to predict objects as out-of-context, resulting in high recall but low accuracy and precision.
-VisSemanticNet achieves the best overall performance.
-Despite their simplicity, our context classification models show promising results and we hope they will inspire further research for more advanced context-aware models.
+Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator.
+5.1 Fine-grained context classification
+Our context classification task addresses a fundamental question: “Does an object belong in this context?” To answer this question effectively at scale, we need models that can perform accurate context reasoning while being practical for deployment.
+While our verification showed that Qwen2.5-VL-72B achieved excellent performance in context classification, its large size makes it impractical for many real-world deployment scenarios.
+To address this challenge, we employ knowledge distillation [36, 24, 63, 53], a proven technique that transfers knowledge from large teacher models to smaller, more efficient student models.
+We use Qwen2.5-VL-72B as the teacher to distill its knowledge into three independent Qwen2.5-VL-3B student models—one for location, one for size, and one for co-occurrence reasoning.
+This distillation process reduces the model size by 24× while maintaining most of the reasoning capabilities, making context classification feasible for practical deployment.
+Note that this knowledge distillation task is possible thanks to the rich context reasoning labels in COinCO.
+We use the 73,929 unanimous consensus images from COinCO to create balanced evaluation splits for training our distilled models.
+Specifically, we sample 24,000 images maintaining a 1:1 ratio of in-context to out-of-context examples, which we divide into 20,000 for training, 2,000 for validation, and 2,000 for testing.
+This balanced sampling ensures unbiased evaluation of context reasoning capabilities.
+We use context reasoning responses from QwenVL as the ground truth (discussed in Section 4.2).
+Importantly, each student model learns not only the final decision (in-context or out-of-context) but also the reasoning process for its specific criterion, enabling interpretable explanations.
+Table 3: Context reasoning accuracy (%) on COinCO test set and original COCO images. Results on original COCO confirm models learn genuine context reasoning rather than inpainting artifacts.
+COinCO Test Set Original COCO Images
+Criterion Before After $\Delta$ Before After $\Delta$
+Size 55.4 65.9 +10.5 52.6 83.2 +30.6
+Location 67.6 75.4 +7.8 47.6 91.4 +43.8
+Co-occurrence 75.5 79.9 +4.4 89.0 87.0 -2.0
+Average 66.2 73.7 +7.6 63.1 87.2 +24.1
+Table 3 shows consistent improvements for each student model across all criteria.
+On the COinCO test set, size reasoning improves most (+10.5% accuracy), while co-occurrence, despite the strongest baseline, reaches nearly 80% accuracy, confirming successful knowledge transfer from the 72B teacher to 3B students.
+Validation on original COCO images.
+To verify that our models learn genuine semantic relationships and context rather than fitting to inpainting artifacts, we test them on 2,000 original COCO images without any manipulation.
+The ground-truth values of all three criteria are true as the objects are considered as in-context.
+The co-occurrence model achieves 89.0% accuracy on original images before fine-tuning, demonstrating it captured natural object relationships.
+After fine-tuning on COinCO, accuracy slightly decreases to 87.0%, suggesting the model already learned robust co-occurrence patterns that need no further adjustment.
+The location model shows the largest improvement after fine-tuning (+43.8% accuracy), indicating that exposure to diverse object placements in COinCO helps the model learn spatial reasoning that transfers to real images.
+The size model also improves substantially (+30.6% accuracy), learning appropriate scale relationships.
+For location and size reasoning, models achieve an average improvement of +37.2% accuracy on original COCO images, with the high performance on unmanipulated images confirming these models understand genuine contextual principles rather than memorizing dataset-specific patterns.
+Each model provides interpretable natural language explanations for its decisions.
+Through distillation, these 3B models achieve practical inference speeds while maintaining reasoning quality [24, 62].
+The modular design allows criterion-specific refinements, demonstrating that complex contextual understanding can be decomposed into focused reasoning tasks.
+While the 72B teacher model is too slow for practical deployment, these efficient 3B student models make context reasoning feasible for real-world applications.
+Refer to caption Refer to caption Refer to caption
+Instance Clique
+skis 31.6 sports 79.2
+snowboard 24.3 accessory 6.6
+kite 14.8 person 4.9
+surfboard 7.0 vehicle 4.2
+backpack 5.9 animal 3.0
+Instance Clique
+kite 22.9 sports 75.1
+skis 22.1 accessory 6.1
+snowboard 19.6 vehicle 5.7
+surfboard 8.6 person 5.5
+person 5.5 animal 4.2
+Instance Clique
+horse 5.0 animal 37.2
+zebra 5.0 vehicle 14.9
+giraffe 4.9 sports 12.1
+person 4.6 outdoor 10.7
+elephant 4.4 accessory 6.8
+Figure 5: Object-from-Context prediction.
+A red box in each image indicates the query region.
+The top row shows three examples (two inpainted, one original COCO).
+The bottom row lists instance-level and clique-level predictions with their probabilities (P(%)).
+Objects in red are the top predictions.
 5.2 Objects-from-Context prediction
-While our context classification model determines if an existing object belongs in a scene, our Objects-from-Context prediction task tackles a different question: “What object(s) fit to this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [52].
+While context classification models determine if an existing object belongs in a scene, our Objects-from-Context prediction tackles a different question: “What object(s) fit this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [55].
 We formulate this task at two levels: 1) Instance-level prediction.
-Predicting the exact object class among COCO’s 80 classes.
+Predicting the object class among COCO’s 80 classes.
 2) Clique-level prediction.
@@
 If no other objects are present, this baseline defaults to random selection.
+We evaluate our Objects-from-Context model on the 2,402 images inpainted from the COCO2017 validation set.
 Table 4: Instance-level and clique-level object-from-context prediction accuracy.
@@
 This result demonstrates that our approach effectively learns the relationship between scene contexts and the objects that belong within them.
-Our model offers three key applications as shown in the three rows of Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [54]; (2) suggesting contextually appropriate objects based on spatial location within a scene, enabling intelligent image editing [33]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation.
+Our model offers three key applications as shown in Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [44]; (2) suggesting contextually appropriate objects based on spatial location within a scene [50, 51], enabling intelligent image editing [11]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation.
 This versatility advances context-driven visual understanding across multiple practical applications.
-Refer to caption Refer to caption Refer to caption
-Instance P(%) skis: 31.6 snowboard: 24.3 kite: 14.8 surfboard: 7.0 backpack: 5.9 Clique P(%) sports: 79.2 accessory: 6.6 person: 4.9 vehicle: 4.2 animal: 3.0 Instance P(%) kite: 22.9 skis: 22.1 snowboard: 19.6 surfboard: 8.6 person: 5.5 Clique P(%) sports: 75.1 accessory: 6.1 vehicle: 5.7 person: 5.5 animal: 4.2 Instance P(%) horse: 5.0 zebra: 5.0 giraffe: 4.9 person: 4.6 elephant: 4.4 Clique P(%) animal: 37.2 vehicle: 14.9 sports: 12.1 outdoor: 10.7 accessory: 6.8
-Figure 5: Object-from-Context prediction.
-A red box is a query.
-The top row shows three examples (two inpainted, one original COCO), and the bottom row lists instance-level and clique-level predictions with their probabilities (P).
-Objects in red are the top predictions.
+5.3 Context-empowered fake localization
+Table 5: SOTA fake detection performance on COinCO.
 Method Acc F1 AUC AP
@@
 CAT-Net [32] 92.7 76.5 97.4 90.3
-Table 5: SOTA fake detection performance.
 Figure 6: Context enhancement results.
-5.3 Context-empowered fake localization
 Fake localization aims to identify the specific synthetic regions in an image, producing a pixel-level map that is fine-grained compared to binary image-level classification.
 In COinCO, the ground truth mask for fake localization is the original object’s entire bounding box used in inpainting.
-We evaluate localization performance using four common pixel-level metrics [41, 40]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
+We evaluate localization performance on the 2,402 test images derived from COCO2017 validation using four common pixel-level metrics [32, 39]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
 State-of-the-art performance.
 We benchmark several SOTA fake detection models on COinCO (Table 5).
-The high accuracy and AUC scores ($>84\%$ for all models) can be attributed to the small average size of inpainted regions in our data.
+The high accuracy and AUC scores ($84\%$ for all models) can be attributed to the small average size of inpainted regions in our data.
 As most images are majority authentic, these metrics are biased towards high values.
@@
 Trufor obtains the second-best F1 score of $55.92\%$, while PSCC-Net shows strong performance in AUC ($95.38\%$) and AP ($79.52\%$).
-ManTraNet demonstrates relatively lower performance, particularly in F1 score ($30.68\%$) and AP ($50.72\%$).
+ManTraNet demonstrates relatively lower performance in F1 score ($30.68\%$) and AP ($50.72\%$).
 These findings highlight significant differences and room for improvement among SOTA fake localizers.
@@
 To leverage contextual information for improved fake localization, we propose a simple yet effective way of enhancing the prediction scores of fake localization models in regions corresponding to out-of-context objects.
-Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; otherwise $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
+Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
 Image GT [32] [60] [20] [39]
@@
 In the practical setting, the fake objects are unknown.
-We propose the use of Molmo to identify suspicious objects based on size and location, deliberately excluding co-occurrence to prevent false positives.
+We propose the use of Molmo-7B [15] to identify suspicious objects based on size and location, excluding co-occurrence to prevent false positives.
 For instance, if an image contains only two objects, e.g., an apple and an inpainted traffic light, either might be seen as out-of-context.
@@
 In both settings, we set the enhancement factor $\gamma$ to $5$ to enhance prediction scores in out-of-context regions.
-A detailed analysis of $\gamma$ is in supplementary materials.
-Figure 6 shows how our context-enhancement improves fake localization for all SOTA detectors without any finetuning.
+The analysis of $\gamma$ is in supplementary materials.
+Figure 6 shows that our context-enhancement improves fake localization for SOTA detectors without any finetuning.
 The effectiveness of our method stems from the complementary nature of context detection and fake localization.
-This combination ensures that even if Molmo misidentifies authentic objects as out-of-context, enhancement only strengthens the base model’s predictions, preserving robustness.
-These results highlight context as a valuable signal for fake detection, opening new possibilities for context-aware image forensics.
-6 Conclusion and Limitations
-We present COinCO, a novel dataset designed to feature in- and out-of-context objects in real scenes.
-By strategically replacing objects in COCO with diffusion-based inpainting, we systematically diversify the contextual status of objects in complex scenes.
-COinCO’s context labels allow us to train a general context classifier.
-We also introduce a novel Object-from-Context prediction task.
-Finally, our work advances current fake detection approaches by leveraging contextual information during fake pixel localization.
-COinCO and our findings highlight the importance of context in visual understanding and provide a foundation and testbed for future research.
-While COinCO advances context-aware visual understanding, context classification remains inherently subjective despite our structured criteria, as human annotators and MLLMs occasionally disagree on contextual coherence.
-The Objects-from-Context prediction task is currently limited to COCO’s predefined categories.
-Future work could explore extending these tasks to open-vocabulary settings for more flexible context reasoning.
+Thus, even if LVLM mistakes authentic objects as out-of-context, enhancement only strengthens the base model’s predictions.
+Notably, out-of-context objects are not rare in real-world forgeries, especially in partial manipulation scenarios, where inserted objects often exhibit semantic inconsistencies in size, location, or scene compatibility.
+Our results highlight that context is valuable for fake detection.
+6 Conclusion
+We present COinCO, a large-scale dataset featuring both in-context and out-of-context objects created through controlled diffusion-based inpainting.
+By strategically replacing COCO objects, COinCO introduces diverse contextual configurations and provides rich criterion-level annotations generated through multiple LVLMs consensus.
+These annotations enable training efficient, interpretable context classifiers through knowledge distillation, and support a new Objects-from-Context prediction task that models what objects naturally belong in a scene.
+We further show that contextual cues serve as a powerful complementary signal for fake localization, improving SOTA detectors without any fine-tuning.
+Overall, COinCO establishes a comprehensive testbed for advancing context-aware visual understanding, contextual reasoning, and image forensics.
+Future work could extend context reasoning to open-vocabulary settings beyond COCO’s predefined categories, enabling more generalizable contextual understanding.
+7 Acknowledgments
+This work was funded in part by the National Science Foundation (SES-2521631) and the University of Georgia’s Presidential Interdisciplinary Seed Grant.